OpenStudy (anonymous):

if xy= -3 and (x+y)^2 = 33, what is the value of x^2 + y^2?

OpenStudy (y2o2):

(x+y)² = x² + 2xy + y² = 33 2xy = -6 so; x²+y² = 39

OpenStudy (anonymous):

you made it seem so easy! thanks!

OpenStudy (y2o2):

It's my pleasure :)